Answer all the questions below The production of
Answer #1

The stoichiometric coefficients:

a = 5.845

b = 0.124

c = 0.062

d = 5.938

e = 1

f = 5

The degree of reduction of the substrate, biomass and product

\gamma s = 4+12-2*6 = 4

\gamma b = 4+20-2*4-3*2=10

\gamma p = 4+7-2*3-3*1=2